# A Systematic Literature Review of Software Engineering Research on Jupyter Notebook

### v1

@@ 5.7.2 Best Practices for Collaborative Use
Researchers showed that it is necessary to develop and validate best practices for collaborative notebook use and the tools required to enforce these practices [119]. Quaranta et al. [121] introduced a catalog of collaboration-specific best practices for Jupyter Notebook, aiming to improve teamwork, reproducibility, and code quality in data science workflows. The catalog emphasizes practices like using version control, structuring code with modular functions, documenting with Markdown, cleaning and organizing cells, and separating exploratory from production notebooks. It also encourages open sharing and the use of self-contained environments to ensure reproducibility. To manage editing conflicts during real-time collaboration effectively, Wang et al. [170] proposed PADLOCK, which provides conflict-resolution mechanisms such as cell-level and variable-level access control and parallel cell groups. Their empirical evaluation demonstrated reduced editing conflicts and improved support for diverse collaboration styles. Together, these guidelines and tools support more maintainable and collaborative notebook development.
Summary of best practices in notebooks Unlike traditional software development environments, Jupyter notebooks often prioritize data processing and exploration, which can lead to deviations from established software engineering practices. No studies have identified or explored the ML-specific best practices in notebooks. Building upon the study by Pimentel et al. [116], Grotov et al. [46], which identified good and bad coding practices, there is a need for an impact analysis of these coding practices.
@@ 5.8 Cell Execution Order
Notebook cells can be executed independently and out of order, making it challenging for developers to manage the global variables that notebooks keep persistent in memory [93]. Developers should strive to maintain a linear order of code execution in Jupyter notebooks, as researchers noticed that the nonlinear execution order is one of the main challenges while reusing notebooks [151, 93]. These non-linear executions in Jupyter notebooks create difficulties for users in understanding and tracing dependencies between code cells, which can cause confusion, errors, and challenges in reproducing results [14]. To address this, Brown et al. [14] introduced Dataflow Notebooks (DFNBs), a system that can locate variable definitions and track variable references as explicit data dependencies to overcome out-of-order cell executions.
@@
Developing reproducible ML pipelines can also help address notebook execution order issues. In this context, the ML pipelines in notebooks consist of interconnected code cells where each step of the ML process (such as preprocessing, model training, and evaluation) is represented as a separate code cell [61, 99]. Jiang et al. [61] developed an approach to determine the correct execution order by identifying and extracting the underlying structure of a notebook by building a labeled dependency graph. In this graph, each cell is represented as a node labeled with a specific ML stage (e.g., data collection, training, evaluation), and edges represent data dependencies between cells. These labeled cells are then reordered to maintain the ML pipeline execution flow. A tool named Ploomber has been developed to manage and automate the execution sequence of notebook cells [99]. Ploomber transforms notebooks into a structured format, making a Jupyter notebook function like a reproducible pipeline with a single execution flow. This tool allows users to break down large notebooks into smaller, manageable tasks that are connected through a clear execution order. This structured approach helps prevent common issues associated with out-of-order execution, such as undefined variables or incorrect data states.
@@
Summary of cell execution order Nonlinear cell execution in Jupyter notebooks leads to persistent memory challenges, making it difficult for developers to track variable dependencies and ensure reproducibility [93, 151]. To address this, researchers have developed methods like DFNBs for tracking dependencies [14], and tools like Ploomber [99] to automate and streamline notebook workflows. Researchers also developed solutions for reproducible ML pipelines, helping resolve notebook execution order issues [61, 99].
@@ 5.9 AI-based Coding Assistance for Notebooks
Jupyter notebooks offer a convenient way to write and execute code in a single shareable file for exploratory data analysis and insight finding [72]. However, users with limited coding experience may struggle to participate in the analysis process quickly. AI-based coding assistance for Jupyter notebooks can be a solution for these users by generating code snippets based on various inputs, such as natural language instructions or notebook contexts. Researchers have explored automatic generation of notebook codes with low-code strategies to support data analysis [22, 201]. For example, Chen et al. [22] developed a low-code interaction panel to recommend follow-up questions to guide the next steps in exploratory data analysis [22]. This approach helps users visualize the structure of their data science workflow through a tree-based representation of the notebook cells. Notebook users can generate code cells based on natural language instructions. Yin et al. [201] proposed an approach to enable users to describe their desired outcome in English and automatically generate the necessary code. However, this solution is designed specifically for tasks involving the Pandas library. Huang et al. [58] further expanded code generation methods by developing DataCoder, a dual-encoder model that generates contextualized data-wrangling code cells by encoding textual, code, and tabular data separately to improve accuracy. Users can also generate notebook code cells by interacting with appropriate UI scaffolds. For instance, Cheng et al. [23] developed BISCUIT, a JupyterLab extension leveraging ephemeral user interfaces generated by LLMs, enabling interactive elements (e.g., sliders, dropdowns) to facilitate more understandable and exploratory ML coding tasks.
@@
Beyond code generation, studies have emphasized the importance of mediated interactions and enhanced code-cell execution assistance. For educational environments, George and Dewan [38] introduced NotebookGPT, embedding GPT interactions within Jupyter notebooks to encourage effective prompt usage and reduce the direct copying behavior among students through programmatic mediation strategies. In parallel, efforts have also been directed toward defining new design approaches for seamlessly integrating LLM support into notebook workflows, potentially improving productivity for diverse developer groups [188]. Similarly, Brault et al. [13] demonstrated a solution for generating new notebooks by retrieving relevant past experiments based on user-specified problem configurations. Lastly, McNutt et al. [97] investigated AI-powered notebook assistants which recommend optimal code cell execution pathways by analyzing data dependencies alongside the code itself, highlighting another crucial dimension of notebook assistance beyond generating code snippets.
@@
Summary of AI-based coding assistance for notebook AI-based coding assistance in Jupyter notebooks helps users generate code snippets using natural language instructions or notebook contexts. Researchers have explored a range of techniques, including low-code interaction panels [22, 201], context-aware AI models [58], and interactive UI scaffolds [23] to streamline notebook code. Additionally, AI-powered assistants integrated into notebooks facilitate code-cell execution [97], workflow guidance [13], and enhanced user interactions [38] to make notebooks more intuitive for broader audiences.
@@ 5.10 Supporting other Programming Paradigms
To support the use of multiple programming languages within a notebook, researchers have proposed polyglot execution environments [104]. A polyglot notebook system allows code cells to interact directly with data structures and invoke functions or methods from different programming languages [104, 114]. This system facilitates direct object sharing and function calling across different programming languages, eliminating the need for data serialization between languages. Hence, a polyglot notebook system can improve the interoperability of various programming languages within a single notebook environment.
Furthermore, block-based visual programming within notebooks provides a graphical interface for programming, making the structure and logic of the code visually clear [168]. Integrating block-based programming into Jupyter notebooks enhances coding practices, simplifies syntax, reduces errors, and lowers the learning curve for non-expert developers. Verano Merino et al. [168] proposed this integration to provide a user-friendly interface for programming and data science tasks, thus facilitating the adoption of computational notebooks across different domains. Similarly, Weber et al. [187] developed a multi-paradigm editor within the Jupyter ecosystem that integrates graphical and textual views with automated synchronization, demonstrating enhanced usability, lower workload, and prevention of execution order errors.
Summary of supporting other programming paradigms Supporting multiple programming paradigms in notebooks enhances flexibility by enabling cross-language execution and intuitive visual programming interfaces. For example, polyglot execution environments enable interaction between multiple programming languages within a notebook [104, 114]. Additionally, block-based visual programming enhances accessibility by providing a graphical interface, simplifying syntax, and supporting non-expert developers in computational notebooks [168, 187].
@@ 5.11 Datasets of Notebooks
2024 https://github.com/ISE-Research/DistilKaggle 542,051 notebooks from Kaggle with 34 code quality metrics [40, 39]
@@
2024 https://doi.org/10.6084/m9.figshare.26372140 297,800 ML notebooks from GitHub and Kaggle to study feedback mechanisms (e.g., assertions and print) [147]
@@
2024 https://zenodo.org/records/11396773 138,376 notebooks from Kaggle used to study data preprocessing practices in ML development [43]
@@
2024 https://github.com/Jun-jie-Huang/CoCoNote 58,221 code generation examples from GitHub Jupyter notebooks to study data-wrangling [58]
@@
2024 https://huggingface.co/datasets/JetBrains-Research/jupyter-errors-dataset 10,000 GitHub notebooks containing at least one thrown exception [47]
@@
2024 https://zenodo.org/records/15114367 64,031 ML notebooks (61k GitHub, 2.7k Kaggle) with 92,542 crashes [182]
@@
2024 https://zenodo.org/records/13836922 113 notebooks containing 342 recommended cells [7]
2022 https://zenodo.org/records/6383115 847,881 notebooks from GitHub to study code structure and style [46]
@@
2022 https://github.com/bugs-jupyter/empirical-study 105 notebooks from GitHub with 14,740 bug-related commits [27]
2021 https://zenodo.org/records/4468523 248,761 notebooks from Kaggle (the KGTorrent dataset) [120, 24, 41, 126, 40, 43]
2021 https://zenodo.org/records/5109482 267,602 notebooks from Kaggle [121]
2021 https://ibm.biz/Bdfpk6 3,944 notebooks from Kaggle [91]
2021 https://zenodo.org/records/7109939 470 notebooks from GitHub for analyzing the workflow of data science code [127]
2020 https://github.com/SMAT-Lab/SnifferDog/blob/master/dataset/all.github.urls 100,000 notebook projects from GitHub, with 507 notebook projects which are executed for reproducibility [180]
@@
2020 https://zenodo.org/records/3836691 6,000 notebooks from GitHub with code clones [73]
2020 https://osf.io/9q4wp 2,574 notebooks from GitHub [125]
2020 https://figshare.com/s/4c5f96bc7d8a8116c271 200 notebooks from GitHub with meaningful commit histories [90]
@@
2019 https://zenodo.org/records/2592524 1.4 million notebooks from GitHub [115, 105, 116, 177, 159]
@@
2018 https://library.ucsd.edu/dc/object/bb2733859v 1.25 million notebooks from GitHub [135, 109, 84]
@@
Publicly available datasets of Jupyter notebooks have been instrumental in advancing research on computational notebooks, enabling large-scale empirical studies on their usage, quality, and evolution. These datasets typically include collections of notebooks from platforms such as GitHub and Kaggle, often annotated with metadata such as commit histories, bug-related information, or workflow analysis. In our systematic literature review, we discovered 18 distinct datasets that are publicly available. Table 7 lists all the public datasets of Jupyter notebooks used in the literature, highlighting their publication years, public URLs, the description of sources, and research articles that used them. Among these, 11 were gathered from GitHub, 5 from Kaggle, and 2 from both GitHub and Kaggle.
@@
These datasets vary significantly in size and purpose. Some datasets are intended for general research on Jupyter Notebook. For instance, the dataset from Grotov et al. [46] contains 847,881 notebooks from GitHub, supporting analyses of notebook usage trends. Similarly, the KGTorrent dataset includes 248,761 notebooks from Kaggle, offering insights into the structured use of notebooks in data science competitions [120]. Additionally, large-scale collections such as the 1.4 million GitHub notebooks compiled by Pimentel et al. [115] and the 1.25 million GitHub notebooks by Rule et al. [135] provide a comprehensive view of the Jupyter Notebook platform. DistilKaggle [39, 40] extends KGTorrent, containing 542,051 Kaggle notebooks annotated with 34 code quality metrics.
@@
Specialized datasets, such as the one by De Santana et al. [26], focus on bug-related commits, including 105 Jupyter notebooks from GitHub with 14,740 associated commits, making it highly relevant for debugging and maintenance studies. Similarly, the dataset from Wang et al. [180] emphasizes the reproducibility of notebooks, while the dataset by Koenzen et al. [73] specifically supports research on code clone detection. Another focused dataset from Ramasamy et al. [127] provides data science code for analyzing the workflow. Furthermore, Grotov et al. [47] built a dataset of 10,000 GitHub notebooks explicitly containing exceptions and error outputs, and Huang et al. [58] focused on the new task of contextualized data-wrangling code generation leveraging code, data, and text contexts. Aydin et al. [7] prepared a CelRec-DB dataset containing 342 recommended code cells worked in Jupyter notebooks. Unfortunately, one of the used datasets [63] was not accessible due to a permission issue and therefore excluded from the table.
@@
Several recent datasets specifically target ML notebooks to address emerging research challenges and trends. For instance, Golendukhina and Felderer [43] curated a dataset of 138,376 Kaggle notebooks that contain ML-specific data preprocessing API usage based on the KGTorrent dataset [120]. Similarly, Wang et al. [182] collected 64,031 ML notebooks that contain execution errors (notebook crashes) to support research on debugging practices and error mitigation. Shome et al. [147] constructed a dataset containing 297,800 ML notebooks from both GitHub and Kaggle, including 89.6 thousand assertions, 1.4 million print statements, and 1 million last cell statements.
@@
Summary of datasets of notebooks Publicly available Jupyter Notebook datasets from GitHub and Kaggle serve various purposes, from analyzing overall notebook usage to studying reproducibility and reusability [180, 73]. Some collections, such as KGTorrent [120] and DistilKaggle [39, 40], focus on structured notebook usage, while others, like CelRec-DB [7], capture recommended code cells. Additionally, some datasets specifically address machine learning notebooks, supporting research on preprocessing API usage [43], execution errors [47], and debugging practices [182, 26].

### v2

@@ 5.7.2 Best Practices for Collaborative Use
Researchers showed that it is necessary to develop and validate best practices for collaborative notebook use and the tools required to enforce these practices [quaranta2022assessing]. quaranta2022eliciting introduced a catalog of collaboration-specific best practices for Jupyter Notebook, aiming to improve teamwork, reproducibility, and code quality in data science workflows. The catalog emphasizes practices like using version control, structuring code with modular functions, documenting with Markdown, cleaning and organizing cells, and separating exploratory from production notebooks. It also encourages open sharing and the use of self-contained environments to ensure reproducibility. To manage editing conflicts during real-time collaboration effectively, wang2024dont proposed PADLOCK, which provides conflict-resolution mechanisms such as cell-level and variable-level access control and parallel cell groups. Their empirical evaluation demonstrated reduced editing conflicts and improved support for diverse collaboration styles. Together, these guidelines and tools support more maintainable and collaborative notebook development.
Summary of best practices in notebooks Unlike traditional software development environments, Jupyter notebooks often prioritize data processing and exploration, which can lead to deviations from established software engineering practices. No studies have identified or explored the ML-specific best practices in notebooks. Building upon the study by pimentel2021understanding, grotov2022large, which identified good and bad coding practices, there is a need for an impact analysis of these coding practices.
@@ 5.8 Cell Execution Order
Notebook cells can be executed independently and out of order, making it challenging for developers to manage the global variables that notebooks keep persistent in memory macke2021automating. Developers should strive to maintain a linear order of code execution in Jupyter notebooks, as researchers noticed that the nonlinear execution order is one of the main challenges while reusing notebooks singer2020notes, macke2021automating. These non-linear executions in Jupyter notebooks create difficulties for users in understanding and tracing dependencies between code cells, which can cause confusion, errors, and challenges in reproducing results [brown2023facilitating]. To address this, brown2023facilitating introduced Dataflow Notebooks (DFNBs), a system that can locate variable definitions and track variable references as explicit data dependencies to overcome out-of-order cell executions. Building on these challenges, in2025exploring explored how immersive computational notebooks can support non-linear execution workflows through spatial organization strategies. Spatial organization in computational notebooks refers to how code cells are visually arranged to reflect their logical relationships and execution order. In a controlled study with 20 participants, in2025exploring found that users preferred half-cylindrical and quarter-cylindrical layouts to manage notebook execution order.
@@
Developing reproducible ML pipelines can also help address notebook execution order issues. In this context, the ML pipelines in notebooks consist of interconnected code cells where each step of the ML process (such as preprocessing, model training, and evaluation) is represented as a separate code cell jiang2022elevating, michael2022keeping. jiang2022elevating developed an approach to determine the correct execution order by identifying and extracting the underlying structure of a notebook by building a labeled dependency graph. In this graph, each cell is represented as a node labeled with a specific ML stage (e.g., data collection, training, evaluation), and edges represent data dependencies between cells. These labeled cells are then reordered to maintain the ML pipeline execution flow. A tool named Ploomber has been developed to manage and automate the execution sequence of notebook cells [michael2022keeping]. Ploomber transforms notebooks into a structured format, making a Jupyter notebook function like a reproducible pipeline with a single execution flow. This tool allows users to break down large notebooks into smaller, manageable tasks that are connected through a clear execution order. This structured approach helps prevent common issues associated with out-of-order execution, such as undefined variables or incorrect data states.
@@
Summary of cell execution order Nonlinear cell execution in Jupyter notebooks leads to persistent memory challenges, making it difficult for developers to track variable dependencies and ensure reproducibility macke2021automating, singer2020notes. To address this, researchers have developed methods like DFNBs for tracking dependencies [brown2023facilitating], and tools like Ploomber [michael2022keeping] to automate and streamline notebook workflows. Researchers also developed solutions for reproducible ML pipelines, helping resolve notebook execution order issues [jiang2022elevating, michael2022keeping].
@@ 5.9 AI-based Coding Assistance for Notebooks
Jupyter notebooks offer a convenient way to write and execute code in a single shareable file for exploratory data analysis and insight finding kluyver2016jupyter. However, users with limited coding experience may struggle to participate in the analysis process quickly. AI-based coding assistance for Jupyter notebooks can be a solution for these users by generating code snippets based on various inputs, such as natural language instructions or notebook contexts. Researchers have explored automatic generation of notebook codes with low-code strategies to support data analysis chen2023whatsnext, yin2023natural. For example, chen2023whatsnext developed a low-code interaction panel to recommend follow-up questions to guide the next steps in exploratory data analysis chen2023whatsnext. This approach helps users visualize the structure of their data science workflow through a tree-based representation of the notebook cells.
Notebook users can generate code cells based on natural language instructions. yin2023natural proposed an approach to enable users to describe their desired outcome in English and automatically generate the necessary code. However, this solution is designed specifically for tasks involving the Pandas library. huang2024contextualized further expanded code generation methods by developing DataCoder, a dual-encoder model that generates contextualized data-wrangling code cells by encoding textual, code, and tabular data separately to improve accuracy. Users can also generate notebook code cells by interacting with appropriate UI scaffolds. For instance, cheng2024biscuit developed BISCUIT, a JupyterLab extension leveraging ephemeral user interfaces generated by LLMs, enabling interactive elements (e.g., sliders, dropdowns) to facilitate more understandable and exploratory ML coding tasks.
AI-assisted code editing in interactive machine learning notebooks faces unique challenges due to their non-linear structure, frequent incremental changes, and rich contextual dependencies. jin2025suggesting, jin2025learning demonstrated that the edits on Jupyter notebooks are highly localized and followed incremental maintenance patterns. AI models are increasingly used to assist with these incremental edits, helping developers maintain and evolve their data science workflows more efficiently [jin2025suggesting, jin2025learning]. Notebook editing has been modeled as a sequence-to-sequence task, where AI systems leverage historical revision data and cell-level context to predict realistic and intent-driven modifications. Despite these advances, state-of-the-art LLMs like DeepSeek-Coder-6.7b-instruct [guo2024deepseek] still struggle to handle the complexity and diversity of real-world ML notebooks [jin2025suggesting].
Notebook users often lose intermediate results during long editing which makes it difficult to reproduce results due to non-linear execution paths. fang2025large demonstrated that AI agents like GPT4o can improve the usability of Jupyter notebooks by automatically checkpointing them after editing during execution. The checkpointing approach helps preserve intermediate states, making it easier to recover from failures and understand the computational history of a notebook. Complementing this, you2025datawiseagent proposed DatawiseAgent, a notebook-centric LLM agent framework that not only integrates execution context and tool usage but also maintains structured agent memory to enhance reproducibility and automate end-to-end data science workflows.
@@
Beyond code generation and modification, studies have emphasized the importance of mediated interactions and enhanced code-cell execution assistance. For educational environments, george2024notebookgpt introduced NotebookGPT, embedding GPT interactions within Jupyter notebooks to encourage effective prompt usage and reduce the direct copying behavior among students through programmatic mediation strategies. Although these solutions need to connect proprietary closed-sourced AI models, wandel2025pyevalai introduced an AI-assisted evaluation system named PyEvalAI, which automatically scores Jupyter notebooks using a combination of unit tests and a locally hosted language model. In parallel, efforts have also been directed toward defining new design approaches for seamlessly integrating LLM support into notebook workflows, potentially improving productivity for diverse developer groups weber2024computational. Similarly, brault2023taming demonstrated a solution for generating new notebooks by retrieving relevant past experiments based on user-specified problem configurations. mcnutt2023design investigated AI-powered notebook assistants that recommend optimal code cell execution pathways by analyzing data dependencies alongside the code itself, highlighting another crucial dimension of notebook assistance beyond generating code snippets.
@@
Summary of AI-based coding assistance for notebook AI-based coding assistance in Jupyter notebooks helps users generate code snippets using natural language instructions or notebook contexts. Researchers have explored a range of techniques, including low-code interaction panels [chen2023whatsnext, yin2023natural], context-aware AI models [huang2024contextualized], and interactive UI scaffolds [cheng2024biscuit] to streamline notebook code. Additionally, AI-powered assistants integrated into notebooks facilitate code-cell execution mcnutt2023design, workflow guidance [brault2023taming], suggesting edits [jin2025suggesting, jin2025learning] and enhanced user interactions [george2024notebookgpt, wandel2025pyevalai] to make notebooks more intuitive for broader audiences.
@@ 5.10 Supporting other Programming Paradigms
To support the use of multiple programming languages within a notebook, researchers have proposed polyglot execution environments [niephaus2019polyjus]. A polyglot notebook system allows code cells to interact directly with data structures and invoke functions or methods from different programming languages niephaus2019polyjus, petricek2018wrattler. This system facilitates direct object sharing and function calling across different programming languages, eliminating the need for data serialization between languages. Hence, a polyglot notebook system can improve the interoperability of various programming languages within a single notebook environment. Extending this line of work, li2025kernel introduced Kernel-FFI, a system that enables transparent foreign function interfaces in notebooks, allowing seamless invocation of functions across language boundaries without requiring manual data marshaling or kernel switching. Kernel-FFI specifically supports object-oriented programming by enabling foreign object referencing and automatic resource management within notebooks.
Furthermore, block-based visual programming within notebooks provides a graphical interface for programming, making the structure and logic of the code visually clear verano2022suppose. Integrating block-based programming into Jupyter notebooks enhances coding practices, simplifies syntax, reduces errors, and lowers the learning curve for non-expert developers. verano2022suppose proposed this integration to provide a user-friendly interface for programming and data science tasks, thus facilitating the adoption of computational notebooks across different domains. Similarly, weber2024extending developed a multi-paradigm editor within the Jupyter ecosystem that integrates graphical and textual views with automated synchronization, demonstrating enhanced usability, lower workload, and prevention of execution order errors.
Summary of supporting other programming paradigms Supporting multiple programming paradigms in notebooks enhances flexibility by enabling cross-language execution and intuitive visual programming interfaces. For example, polyglot execution environments enable interaction between multiple programming languages within a notebook niephaus2019polyjus, petricek2018wrattler. Additionally, block-based visual programming enhances accessibility by providing a graphical interface, simplifying syntax, and supporting non-expert developers in computational notebooks verano2022suppose, weber2024extending.
@@ 5.11 Datasets of Notebooks
2025 https://doi.org/10.5281/zenodo.14281690 48,398 notebook edits derived from 792 Github ML repositories [jin2025suggesting]
2025 https://zenodo.org/records/13357570 21 notebooks with a log of programmers’ activities in Jupyter Notebook [nakamaru2025jupyter]
@@
2024 https://github.com/ISE-Research/DistilKaggle 542,051 notebooks from Kaggle with 34 code quality metrics [mostafavi2024distilkaggle, ghahfarokhi2024predicting]
@@
2024 https://doi.org/10.6084/m9.figshare.26372140 297,800 ML notebooks from GitHub and Kaggle to study feedback mechanisms (e.g., assertions and print) [shome2024understanding]
@@
2024 https://zenodo.org/records/11396773 138,376 notebooks from Kaggle used to study data preprocessing practices in ML development [golendukhina2024unveiling]
@@
2024 https://github.com/Jun-jie-Huang/CoCoNote 58,221 code generation examples from GitHub Jupyter notebooks to study data-wrangling [huang2024contextualized]
@@
2024 https://huggingface.co/datasets/JetBrains-Research/jupyter-errors-dataset 10,000 GitHub notebooks containing at least one thrown exception [grotov2024untangling]
@@
2024 https://zenodo.org/records/15114367 64,031 ML notebooks (61k GitHub, 2.7k Kaggle) with 92,542 crashes [wang2024why]
@@
2024 https://zenodo.org/records/13836922 113 notebooks containing 342 recommended cells [aydin2024cellrecommend]
@@
2024 https://github.com/bugs-jupyter/empirical-study 105 notebooks from GitHub with 14,740 bug-related commits [de2024bug]
@@
2022 https://zenodo.org/records/6383115 847,881 notebooks from GitHub to study code structure and style [grotov2022large]
2021 https://zenodo.org/records/4468523 248,761 notebooks from Kaggle (the KGTorrent dataset) [quaranta2021kgtorrent, choetkiertikul2023mining, mostafavi2024beyond, ragkhitwetsagul2024typhon, mostafavi2024distilkaggle, golendukhina2024unveiling]
@@
2021 https://zenodo.org/records/5109482 267,602 notebooks from Kaggle [quaranta2022eliciting]
2021 https://ibm.biz/Bdfpk6 3,944 notebooks from Kaggle [liu2021haconvgnn]
2021 https://zenodo.org/records/7109939 470 notebooks from GitHub for analyzing the workflow of data science code [ramasamy2023visualising]
2020 https://github.com/SMAT-Lab/SnifferDog/blob/master/dataset/all.github.urls 100,000 notebook projects from GitHub, with 507 notebook projects which are executed for reproducibility [wang2021restoring]
2020 https://zenodo.org/records/3836691 6,000 notebooks from GitHub with code clones [koenzen2020code]
2020 https://osf.io/9q4wp 2,574 notebooks from GitHub [raghunandan2022code]
2020 https://figshare.com/s/4c5f96bc7d8a8116c271 200 notebooks from GitHub with meaningful commit histories [liu2023refactoring]
2019 https://zenodo.org/records/2592524 1.4 million notebooks from GitHub [pimentel2019large, oli2021automated, pimentel2021understanding, wang2020assessing, titov2022resplit]
@@
2018 https://library.ucsd.edu/dc/object/bb2733859v 1.25 million notebooks from GitHub [rule2018exploration, patra2022nalin, li2021nbsearch]
@@
Publicly available datasets of Jupyter notebooks have been instrumental in advancing research on computational notebooks, enabling large-scale empirical studies on their usage, quality, and evolution. These datasets typically include collections of notebooks from platforms such as GitHub and Kaggle, often annotated with metadata such as commit histories, bug-related information, or workflow analysis. In our systematic literature review, we discovered 21 distinct datasets that are publicly available. Table 8 lists all the public datasets of Jupyter notebooks used in the literature, highlighting their publication years, public URLs, the description of sources, and research articles that used them. Among these, 12 were gathered from GitHub, 5 from Kaggle, 2 from both GitHub and Kaggle, and the rest 2 did not mention their data source.
@@
These datasets vary significantly in size and purpose. Some datasets are intended for general research on Jupyter Notebook. For instance, the dataset from grotov2022large contains 847,881 notebooks from GitHub, supporting analyses of notebook usage trends. Similarly, the KGTorrent dataset includes 248,761 notebooks from Kaggle, offering insights into the structured use of notebooks in data science competitions quaranta2021kgtorrent. Additionally, large-scale collections such as the 1.4 million GitHub notebooks compiled by pimentel2019large and the 1.25 million GitHub notebooks by rule2018exploration provide a comprehensive view of the Jupyter Notebook platform. DistilKaggle [ghahfarokhi2024predicting, mostafavi2024distilkaggle] extends KGTorrent, containing 542,051 Kaggle notebooks annotated with 34 code quality metrics.
@@
Specialized datasets, such as the one by de2024bug, focus on bug-related commits, including 105 Jupyter notebooks from GitHub with 14,740 associated commits, making it highly relevant for debugging and maintenance studies. Similarly, the dataset from wang2021restoring emphasizes the reproducibility of notebooks, while the dataset by koenzen2020code specifically supports research on code clone detection. Another focused dataset from ramasamy2023visualising provides data science code for analyzing the workflow. Furthermore, grotov2024untangling built a dataset of 10,000 GitHub notebooks explicitly containing exceptions and error outputs, and huang2024contextualized focused on the new task of contextualized data-wrangling code generation leveraging code, data, and text contexts. aydin2024cellrecommend introduced CelRec-DB dataset, which consists of structured representations of pre-processed Python code cells encoded as semantic vector embeddings (generated using a pre-trained CodeBERT model). Each cell is accompanied by an automatically assigned task label indicating the type of machine learning activity, such as data preprocessing or model training. To understand data science workflows in Jupyter Notebook environments, another dataset has been created with a detailed log of data science programmers’ activities [nakamaru2025jupyter]. Unfortunately, one of the used datasets kallen2021jupyter was not accessible due to a permission issue and therefore excluded from the table.
@@
Several recent datasets specifically target ML notebooks to address emerging research challenges and trends. For instance, golendukhina2024unveiling curated a dataset of 138,376 Kaggle notebooks that contain ML-specific data preprocessing API usage based on the KGTorrent dataset [quaranta2021kgtorrent]. Similarly, wang2024why collected 64,031 ML notebooks that contain execution errors (notebook crashes) to support research on debugging practices and error mitigation. shome2024understanding constructed a dataset containing 297,800 ML notebooks from both GitHub and Kaggle, including 89.6 thousand assertions, 1.4 million print statements, and 1 million last cell statements. Lastly, jin2025suggesting presented 48,398 notebook edits derived from 792 GitHub ML repositories to understand the modification of ML code in notebooks.
@@
Summary of datasets of notebooks Publicly available Jupyter Notebook datasets from GitHub and Kaggle serve various purposes, from analyzing overall notebook usage to studying reproducibility and reusability [wang2021restoring, koenzen2020code]. Some collections, such as KGTorrent [quaranta2021kgtorrent] and DistilKaggle [ghahfarokhi2024predicting, mostafavi2024distilkaggle], focus on structured notebook usage, while others, like CelRec-DB [aydin2024cellrecommend], capture recommended code cells. Additionally, some datasets specifically address machine learning notebooks, supporting research on preprocessing API usage [golendukhina2024unveiling], execution errors [grotov2024untangling], and debugging practices [wang2024why, de2024bug].